# SAM: A Stability-Aware Cache Manager for Multi-Tenant Embedded Databases

### v1

@@ 6. Experimental Evaluation
This section presents a comprehensive experimental evaluation of our proposed system, SAM, which is powered by our novel AURA allocation algorithm. We first detail the experimental setup and baselines, then present a series of targeted experiments designed to validate SAM’s performance, stability, and scalability.
@@ 6.1. General Experimental Setup
System Environment. All experiments run on Windows 11 (Intel Core i5-12600KF, 32GB RAM) with Python 3.11.7 and APSW/SQLite 3.40.0. We stress I/O by using large records (2KB).
@@
We omit Reinforcement Learning (RL) baselines as our target non-stationary workloads and resource-constrained edge nodes present significant challenges for the sample efficiency and continuous retraining requirements of modern RL agents. Furthermore, our Online Frank-Wolfe based SAM-Core already provides a strong theoretical regret bound, making complex RL-based approaches orthogonal to this work’s core contributions.(Mao et al., 2016, 2019)
Metrics. We evaluate both instantaneous performance and sustained performance stability through the following metrics: Scalability (CPU decision time), Productivity (total TPS, P99 latency), Efficiency (hit rate, TPS/Page), and Decision Stability (throughput stability $\sigma_{\text{TPS}}$), and Fairness (QoS for non-hotspot tenants).
Figure 5. S0’s robustness against cache pollution. (a) The attack slightly raises the ‘ops*miss_rate’ score; (c) B7 catastrophically reallocates resources, causing (b) severe VIP degradation. S0 maintains protection.
@@ 6.2. Performance on Standard Dynamic Workloads
Strategy Median TPS Adaptation Lag (s) Allocation Stability ($\sigma$) Decision Overhead
SAM (ours) 34.0 38.7 17.5 O(K) 0.13ms
B11 (Regression) 33.8 53.0 41.4 O(K) ¡0.01ms
B13 (UCP) 32.2 55.6 21.8 O(P·K) 0.02ms
@@
Phase 2 (High-Priority Hotspot)
Median TPS 33.2 19.0 33.5
@@
Phase 3 (Hotspot Shift)
Median TPS 35.2 21.8 35.0
@@
Comparison with Near-Optimal Baselines. To precisely anchor SAM’s performance, we compare it against an idealized global LRU pool (B5) and the theoretical Hindsight Optimal policy (B14). As summarized in Table 3, the results conclusively show that SAM operates at the theoretical performance limit. Its throughput and hit rate are nearly identical to the optimal B14 baseline in both phases, while decisively outperforming the architecturally-idealized B5 policy. Crucially, SAM’s priority-aware design translates to superior Quality of Service, achieving the lowest P99 latency (77.5 ms) for the high-priority tenant—even surpassing the purely hit-rate-focused oracle policy. This establishes SAM as a near-optimal and practically superior solution.
@@ 6.3. Robustness under Adversarial Workloads
Having established the strong performance of both SAM (S0) and the greedy heuristic (B7) on a standard dynamic workload, we now subject them to a more challenging, adversarial scenario. To validate resilience against (C2) cache pollution, we designed an intermittent scan attack where a low-priority tenant mimics a high-demand workload through cache-unfriendly, scan-intensive operations. This stress-test is designed to answer a critical question: can the policy distinguish between a legitimate hotspot and a deceptive “polluter”? The full failure cascade for the reactive policy is visualized in Figure 5.
Figure 6. Performance comparison of S0 and its ablation variants. The results demonstrate that removing the $\mathcal{H}$-factor (B10) or the $\mathcal{V}$-factor (B8), or making the $\mathcal{H}$-factor overly reactive (B9), all degrade overall performance. The apparent superiority of B3 is a QoS trade-off analyzed in Figure 7.
Figure 7. Quality of Service (QoS) comparison in the complex Phase 3 workload. The figure reveals the critical trade-off where B3’s “winner-take-all” strategy significantly harms non-hotspot tenants, unlike S0’s QoS protection mechanism.
Figure 8. Cache allocation dynamics of the ablation variants. The plot demonstrates that the $\mathcal{V}$-factor is critical for agility (cf. B8’s sluggishness), the $\mathcal{H}$-factor is essential for action (cf. B10’s inertness), and that the $\mathcal{H}$-factor must be conservative (cf. B9’s suboptimal and unstable behavior).
Flawed Logic and Severe Misallocation. An analysis of the resource allocation logic reveals the fundamental weakness of the shallow heuristic. The intermittent scan attack generated a mere 11.7% increase in the ‘ops*miss_rate’ “need score” for the attacker. In response, B7 irrationally increased its cache allocation to this low-value tenant by a staggering 64.5%—a quantified 5.5$\times$ overreaction. To evaluate the consequences of this flawed logic, we created a scenario where a high-priority “VIP” tenant and the attacker compete for a shared 2MB cache. SAM correctly prioritized the VIP, allocating resources at a healthy 1.4-to-1 ratio in its favor (241 vs. 172 pages). In stark contrast, B7’s logic led to a severe resource inversion, diverting the vast majority of resources to the attacker at a disastrous 1-to-4.6 ratio (a mere 73 pages for the VIP vs. 336 for the attacker).
Performance Impact and Final Verdict. This severe misallocation translated directly into degraded service quality. As shown in the time-series plot, SAM maintained the VIP tenant’s hit rate between 97-98% throughout the attack. In contrast, B7’s VIP hit rate fell from 97% to 95.5% during the sustained attack, with a sharp dip to 92%. This experiment demonstrates that B7’s simple heuristic, while effective on benign workloads, becomes a critical vulnerability under adversarial conditions. SAM’s dual-factor design—evaluating both historical contribution and marginal gain—provides the robustness necessary for sustained high performance in real-world systems where such attacks are inevitable.
@@ 6.4. Ablation Studies
To justify the design choices of our SAM algorithm that considers multiple orthogonal factors, we conducted a series of ablation studies. We evaluate the full SAM policy (henceforth S0 in figures) against four variants, each with a critical component modified or removed: B10 (no $\mathcal{H}$-factor), B8 (no $\mathcal{V}$-factor), B3 (no Fixed Pool), and B9 (a fast, reactive $\mathcal{H}$-factor). The results are shown in Figures 6, 7, and 8.
Essentiality and Nature of the $\mathcal{H}$ and $\mathcal{V}$ Factors. The results unequivocally demonstrate that both the $\mathcal{H}$ and $\mathcal{V}$ factors are essential, and that the $\mathcal{H}$-factor must be deliberately conservative. As shown in the overall performance summary (Figure 6a), removing either factor is detrimental. B10 (no $\mathcal{H}$-factor) is the worst performer with a median TPS of 32.4, while B8 (no $\mathcal{V}$-factor) also performs poorly at 33.0 TPS. The allocation dynamics in Figure 8 reveal the reasons for these failures. B10 is completely inert, failing to identify any valuable tenant. Conversely, B8 is sluggish; without the forward-looking $\mathcal{V}$-factor, it is slow to adapt to the new hotspot in Phase 2.
The case of B9, which uses a fast EMA for its $\mathcal{H}$-factor, provides a deeper insight. Despite being more reactive to changes, its overall performance (32.7 TPS) is worse than the more conservative B8. The reason is revealed in its dynamics (Figure 8): its fast-rising $\mathcal{H}$-factor causes it to prematurely converge on a suboptimal allocation of only  300 pages in Phase 2. Furthermore, at the start of Phase 3, this lingering (but quickly-formed) historical bias causes it to mistakenly allocate resources to the now-cold high-priority database. This proves that the $\mathcal{H}$-factor’s role as a stable, conservative anchor (using a slow EMA) is critical, not to impede adaptation, but to ensure consistent, high-quality decisions that lead to sustained performance.
The Value of the QoS Protection Mechanism. The most nuanced result comes from comparing S0 with B3 (no Fixed Pool). While B3 achieves a slightly higher median TPS (34.2 vs. S0’s 34.1, Figure 6a), this minor gain masks a severe trade-off in service quality. Figure 7 provides the crucial insight. In the more complex Phase 3, to achieve a marginal +0.8% gain in hotspot hit rate (92.3% vs. S0’s 91.5%), B3 allows the median non-hotspot hit rate to drop to 96.9%. S0, in contrast, successfully protects these tenants, maintaining their hit rate at 97.9%. This represents a nearly 50% relative increase in cache misses for all other services in the system, demonstrating that B3’s “winner-take-all” strategy is detrimental to overall system health.
Summary. The ablation studies confirm that SAM’s design is a necessary synthesis of its components. The results prove three core design principles: (1) The $\mathcal{H}$-$\mathcal{V}$ dual-factor system is essential for achieving both decision stability and adaptiveness. (2) The $\mathcal{H}$-factor must be conservative and stable to properly anchor the policy’s exploration. (3) The QoS protection mechanism is critical for providing system-wide fairness in a multi-tenant environment.
@@ 6.5. Scalability and Architectural Analysis
Figure 9. Scalability analysis of SAM. The top plot shows SAM’s near-constant decision latency, which outperforms the linear scaling of the greedy heuristic (B7) at larger scales. The bottom plot reveals the reason: SAM’s two-state architectural signature, consisting of a dominant peak of ultra-fast localized optimizations and a sparse tail of rare, more expensive global scans.
@@
6.6. Empirical Checks for Theoretical Guarantees
This subsection empirically validates the key theoretical claims made in §5. We specifically aim to demonstrate: (i) that SAM-Core achieves the optimal $\mathcal{O}(\sqrt{T})$ static regret bound for an online algorithm; and (ii) that the full SAM algorithm’s stability mechanisms successfully bound the allocation change, leading to a regret that is asymptotically identical to its theoretical core.
Method. We rerun the 10-tenant workload (Phases 1–3) and log per-round loss $f_{t}$, oracle loss $f_{t}^{\star}$, and allocation change $\Delta_{t}$. Regret is plotted on a log–log scale; jitter is summarised by the standard deviation $\sigma_{\!\Delta}$ over the steady-state window (last 10% rounds). Each curve averages 5 runs; bands show 95% CI.
@@ 7. Related Work
Resource Management in Multi-Tenant and Cloud Environments. Managing shared resources to provide performance isolation is a central challenge in multi-tenant systems. Seminal work like SQLVM (Narasayya et al., 2013) and large-scale cloud databases like Amazon Aurora (Verbitski et al., 2017) have developed techniques for resource management in monolithic, white-box architectures. Other research has focused on mitigating the “noisy neighbor” problem for CPU (Das et al., 2013) and I/O (Gulati et al., 2010).
In stark contrast, SAM is specifically designed for the unique challenges of edge computing environments, where co-located services are often fully independent, heterogeneous, black-box processes. Our framework acts as an external, non-intrusive coordinator, providing intelligent resource management without requiring any modification to the database engines themselves. The design philosophy of our Two-Pool Architecture is inspired by mature paradigms from system-level resource management like Kubernetes (Burns and Oppenheimer, 2016) resource ‘requests’ and ‘limits’, and memory ballooning in hypervisors (Waldspurger, 2002).
Online Learning and Control for Resource Management. Our work builds upon Online Convex Optimization (OCO) (Hazan, 2016) for theoretical foundations and relates to utility-based cache partitioning like UCP (Qureshi and Patt, 2006). However, SAM advances this paradigm with a novel dual-factor model synthesizing historical efficiency ($\mathcal{H}$-factor) and marginal gain ($\mathcal{V}$-factor), providing a robust framework for the exploitation-exploration trade-off (Bubeck et al., 2012). Our stability mechanisms draw from control theory (Åström and Murray, 2021). SAM’s novelty lies in recognizing that stability is not a constraint but an enabler of sustained performance, and in the synergistic synthesis of online learning and control theory principles to achieve this stability-aware design.
@@ 8. Conclusion
In this paper, we introduced SAM, a stability-aware cache manager that addresses the critical challenge of resource contention in multi-tenant embedded databases. Our key insight is that sustained high performance requires principled stability, not reactive adaptation. SAM’s core, the novel AURA algorithm, implements this philosophy by balancing historical efficiency ($\mathcal{H}$-factor) and forward-looking marginal gain ($\mathcal{V}$-factor). Built on a Two-Pool model to decouple guarantees from optimization, SAM proved superior in extensive evaluation against 14 baselines, achieving near-optimal throughput while uniquely resisting adversarial workloads. These results validate our thesis: decision stability is the foundation for sustained high performance in real-world systems. Future work could extend these principles to proactive allocation, energy efficiency, or multi-resource co-management.

### v2

@@ 6. Experimental Evaluation
This section presents a comprehensive experimental evaluation of our proposed system, SAM, which is governed by its principled AURA control policy. We first detail the experimental setup and baselines, then present a series of targeted experiments designed to validate SAM’s performance, stability, and scalability.
@@ 6.1. General Experimental Setup
Our evaluation is designed to validate SAM’s performance, robustness, and scalability in a controlled, reproducible single-node environment that intentionally stresses memory contention.
System Environment. All experiments were conducted in a controlled local area network environment. The database instances run on a dedicated Windows 11 machine (Intel Core i5-12600KF, 32GB RAM) within a WSL2 environment, while the SAM coordinator and workload generators run on a separate WSL2 instance on the same LAN. This setup is designed to simulate a realistic edge deployment where the coordinator manages co-located databases over a low-latency network connection. To create a challenging, memory-bound environment, we used large record sizes (2KB) and high concurrency (16 threads per instance). Our implementation is based on Python 3.11.7 and SQLite 3.40.0.
@@
We omit Reinforcement Learning (RL) baselines as our target non-stationary workloads and resource-constrained edge nodes present significant challenges for the sample efficiency and continuous retraining requirements of modern RL agents. Furthermore, our Online Frank-Wolfe based AURA-Core already provides a strong theoretical regret bound, making complex RL-based approaches orthogonal to this work’s core contributions (Mao et al., 2016, 2019).
Metrics. We evaluate each strategy along four key dimensions: Scalability (CPU decision time), Productivity (total TPS, P99 latency), Efficiency (hit rate), and Decision Stability. The latter is quantified as the throughput stability ($\sigma_{TPS}$), which we define as the standard deviation of TPS measured over a sliding window during the steady-state period of each experimental phase.
Discussion on Hyperparameter Philosophy. The key hyperparameters in AURA, such as EMA smoothing factors and the momentum weight, do not present a classic optimization problem with a single ‘optimal’ solution. Instead, they define the fundamental ‘personality’ of the system—its position on the spectrum between aggressive responsiveness and conservative stability. An aggressive, fast-reacting configuration might excel under highly dynamic workloads but risk instability from transient noise. Conversely, a conservative, slow-moving configuration ensures deep stability but may adapt too slowly to genuine workload shifts.
Recognizing that no single static personality is universally optimal, AURA’s core design is built to adapt at a higher level. Rather than relying on fine-tuning these fixed hyperparameters, it uses the meta-adaptive $\alpha$ policy to dynamically arbitrate between exploitation ($\mathcal{H}$-factor) and exploration ($\mathcal{V}$-factor) based on perceived environmental uncertainty. Therefore, our evaluations do not aim to prove the universal optimality of our default parameters. Instead, they demonstrate that this chosen ‘balanced personality’ provides robust, high performance across diverse standard and adversarial conditions, validating the effectiveness of this higher-level adaptive design.
@@ 6.2. Performance on Standard Dynamic Workloads
Strategy TPS Lag (s) Stability ($\sigma$) Overhead
SAM (ours) 34.0 38.7 17.5 0.13ms ($O(K)$)
B11 (Regression) 33.8 53.0 41.4 ¡0.01ms ($O(K)$)
B13 (UCP) 32.2 55.6 21.8 0.02ms ($O(P{\cdot}K)$)
@@
Phase 2: High-Priority Hotspot
TPS (Median) 33.2 19.0 33.5
@@
Phase 3: Hotspot Shift
TPS (Median) 35.2 21.8 35.0
@@
Comparison with Near-Optimal Baselines. To precisely anchor SAM’s performance, we compare it against an idealized global LRU pool (B5) and the theoretical Hindsight Optimal policy (B14). The results in Table 3 conclusively show that SAM operates at the theoretical performance limit, with throughput and hit rates nearly identical to the oracle policy. Crucially, SAM’s priority-aware design achieves the lowest P99 latency for the high-priority tenant, even surpassing the purely hit-rate-focused oracle, establishing it as a near-optimal and practically superior solution.
@@ 6.3. Robustness under Adversarial Workloads
We now test SAM’s resilience against cache pollution (C2) by subjecting it and the best greedy heuristic (B7) to an adversarial, intermittent scan attack. This stress-test probes whether a policy can distinguish a legitimate hotspot from a deceptive “polluter”, with the full failure cascade for the reactive policy visualized in Figure 5.
Figure 5. S0’s robustness against cache pollution. (a) The attack slightly raises the ‘ops×$\times$×miss_rate’ score; (c) B7 catastrophically reallocates resources, causing (b) severe VIP degradation. S0 maintains protection.
Figure 6. Ablation studies for SAM (S0). The results confirm that removing the $\mathcal{H}$-factor (B10), the $\mathcal{V}$-factor (B8), or using a reactive $\mathcal{H}$-factor (B9) all degrade performance, while the higher throughput of B3 (no fixed pool) comes at a significant QoS cost (see Figure 7).
Flawed Logic and Severe Misallocation. An analysis of the resource allocation logic reveals the fundamental weakness of the shallow heuristic. The intermittent scan attack generated a mere 11.7% increase in the ‘ops×$\times$×miss_rate’ “need score” for the attacker. In response, B7 irrationally increased its cache allocation to this low-value tenant by a staggering 64.5%—a quantified 5.5$\times$ overreaction. To evaluate the consequences of this flawed logic, we created a scenario where a high-priority “VIP” tenant and the attacker compete for a shared 2MB cache. SAM correctly prioritized the VIP, allocating resources at a healthy 1.4-to-1 ratio in its favor (241 vs. 172 pages). In stark contrast, B7’s logic led to a severe resource inversion, diverting the vast majority of resources to the attacker at a disastrous 1-to-4.6 ratio (a mere 73 pages for the VIP vs. 336 for the attacker).
Performance Impact and Final Verdict. This severe misallocation translated directly into degraded service quality. As shown in the time-series plot, SAM maintained the VIP tenant’s hit rate between 97-98% throughout the attack. In contrast, B7’s VIP hit rate fell from 97% to 95.5% during the sustained attack, with a sharp dip to 92%. This experiment demonstrates that B7’s simple heuristic, while effective on benign workloads, becomes a critical vulnerability under adversarial conditions. SAM’s dual-factor design provides the robustness necessary for sustained high performance in real-world systems where such attacks are inevitable.
@@ 6.4. Ablation Studies
Figure 7. QoS analysis in Phase 3. SAM (S0)’s fixed pool successfully protects non-hotspot tenants, whereas the purely elastic B3 sacrifices their performance for a marginal gain on the hotspot.
Figure 8. Allocation dynamics of the ablation variants, confirming the $\mathcal{V}$-factor’s role in agility (vs. B8’s sluggishness), the $\mathcal{H}$-factor’s role in driving allocation (vs. B10’s inertness), and the necessity of a conservative $\mathcal{H}$-factor (vs. B9’s instability).
To justify SAM’s multi-faceted design, we conducted ablation studies against four variants: B10 (no $\mathcal{H}$-factor), B8 (no $\mathcal{V}$-factor), B3 (no Fixed Pool), and B9 (a reactive $\mathcal{H}$-factor). The results (Figs. 6-8) unequivocally prove that SAM’s components are a necessary synthesis. Removing either the $\mathcal{H}$-factor (B10) or the $\mathcal{V}$-factor (B8) is detrimental, resulting in policy inertness and sluggish adaptation, respectively. The B9 variant, which uses a fast EMA for the $\mathcal{H}$-factor, performs even worse than the conservative B8, proving that the $\mathcal{H}$-factor’s role as a stable, historical anchor is critical to temper the $\mathcal{V}$-factor’s exploratory nature and ensure high-quality decisions. Finally, while B3 achieves a marginally higher throughput, this gain masks a severe QoS trade-off; it allows a nearly 50% relative increase in cache misses for non-hotspot tenants. This confirms that the two-pool architecture is essential for providing system-wide fairness and preventing a ”winner-take-all” strategy that is detrimental to overall system health.
@@ 6.5. Scalability and Architectural Analysis
Figure 10. Scalability analysis of SAM. Its near-constant decision latency outperforms the linear scaling of B7 at scale, a direct result of its two-state architectural signature revealed in the probability density: a dominant peak of ultra-fast localized optimizations and a sparse tail of rare global scans.
@@
6.6. Empirical Validation of Theoretical Guarantees
@@ 7. Related Work
Resource Management in Multi-Tenant and Cloud Environments. Managing shared resources to provide performance isolation is a central challenge in multi-tenant systems. Seminal work like SQLVM (Narasayya et al., 2013) and large-scale cloud databases like Amazon Aurora (Verbitski et al., 2017) have developed techniques for monolithic, white-box architectures, while other research has focused on mitigating the ”noisy neighbor” problem for CPU (Das et al., 2013) and I/O (Gulati et al., 2010). In stark contrast, SAM targets edge computing environments where co-located services are independent, heterogeneous, black-box processes. Acting as an external, non-intrusive coordinator, SAM’s design philosophy draws from mature system-level paradigms like Kubernetes (Burns and Oppenheimer, 2016) resource ‘requests’ and ‘limits’, and memory ballooning in hypervisors (Waldspurger, 2002), positioning it as a distinct alternative to industrial solutions like Facebook’s CacheLib, which optimizes within a single cache instance rather than coordinating across multiple black-box databases.
Online Learning and Control for Resource Management. Our work builds upon Online Convex Optimization (OCO) (Hazan, 2016) for theoretical foundations and relates to utility-based cache partitioning like UCP (Qureshi and Patt, 2006). SAM applies these established principles through a practical dual-factor approach that balances historical efficiency ($\mathcal{H}$-factor) and marginal gain ($\mathcal{V}$-factor), addressing the classic exploitation-exploration trade-off (Bubeck et al., 2012). Our stability mechanisms draw from control theory (Åström and Murray, 2021). SAM’s contribution is in demonstrating that stability-first design—treating stability as an enabler rather than a constraint—can achieve sustained high performance through careful engineering of known techniques.
@@ 8. Conclusion
In this paper, we introduced SAM, a stability-aware cache manager that addresses the critical challenge of resource contention in multi-tenant embedded databases. Our key insight is that sustained high performance requires principled stability, not reactive adaptation. SAM implements this philosophy through its AURA control policy, which balances historical efficiency ($\mathcal{H}$-factor) and forward-looking marginal gain ($\mathcal{V}$-factor). Built on a Two-Pool model to decouple guarantees from optimization, SAM proved superior in extensive evaluation against 14 baselines, achieving near-optimal throughput while uniquely resisting adversarial workloads. These results validate our thesis: decision stability is the foundation for sustained high performance in real-world systems. Future work could extend these principles to proactive allocation, energy efficiency, or multi-resource co-management.
Limitations and Future Work. While SAM demonstrates significant advancements, our work also opens several promising research avenues. Our current model does not explicitly account for the transient costs of cache warming, integrating a cost model for reallocations is a key next step. Another direction is to evolve our reactive strategy into a proactive allocation framework by incorporating lightweight workload prediction. Finally, the principles developed in SAM could be generalized to create a holistic framework for multi-resource co-management, coordinating not only cache but also other contended resources like I/O bandwidth and CPU priority, a critical challenge for future edge computing platforms.